OpenStudy (anonymous):

Number of orbitals that could be associated with 'n' are; a) n^2 + 2 b) 2n^2 c) n^2 d) n^2-2

OpenStudy (anonymous):

The answer is b) 2n^2 If n = 1, then there only exists the S level, which contains 2 orbitals. so 2 * (1^2) = 2 If n = 2, there is the S level and P level, which contains 2 orbitals + 6 orbitals. so 2 * (2^2) = 8. based on these relationships, 2n^2 shows how many orbitals there are according to the value of 'n'

OpenStudy (mos1635):

orbitals refers to elektrons???

OpenStudy (mos1635):

if not then korect ans is n^2

OpenStudy (anonymous):

I am very sorry! My above post was incorrect. I was under the impression that we were talking about the number of electrons in each "n" level. I wrote "orbitals" in the place of what should be "electrons." mos1635 is correct. The relationship between the number of orbitals and "n" is n^2

OpenStudy (anonymous):

So the correct answer is C) n^2

OpenStudy (anonymous):

ok,... thanks :)